### v6

# Dual-Space Invariance as a Universal Criterion for Multifractal Critical States

Abstract
In Anderson localization, eigenstates of disordered quantum systems are broadly classified as extended, localized, or critical. Although critical states exhibit multifractal character, a precise and operational criterion for their identification remains an open challenge, as Lyapunov exponents in real space cannot uniquely distinguish them from extended states. Here we address this challenge by asserting that critical states are uniquely characterized by an emergent dual-space invariance between position and momentum space. Building on the Liu–Xia criterion of the simultaneous vanishing of Lyapunov exponents ($\gamma=\gamma_{m}=0$), we show that this dual-space invariance extends beyond Lyapunov exponents and governs wavefunction scaling, revealing a fundamental property inaccessible from either space alone. Through numerical simulations, we demonstrate that the inverse participation ratio exhibits matching scaling behavior in position and momentum space for critical states, in sharp contrast to extended and localized states, which display a pronounced asymmetry between the two spaces. This dual-space invariance provides a direct, robust, and universal criterion for identifying multifractal critical states. Our results establish a fundamental principle of Anderson criticality and open new avenues for its detection in modern quantum simulation platforms.
I Introduction
Anderson localization photonic ; Critical0 is a fundamental manifestation of wave interference in disordered media, leading to the suppression of diffusive transport Liu-self-duality ; Li1 ; Li2 . Within such systems, single-particle eigenstates are commonly classified as extended (metallic), localized (insulating), or critical. Critical states–also referred to as multifractal states–occupy an intermediate regime between extended and localized phases You ; g1 ; zhou . Unlike extended states, which exhibit nearly uniform probability distributions, or localized states, which decay exponentially, critical states typically display power-law spatial decay Zhang ; q2 . Their hallmark features include multifractality, characterized by a continuous spectrum of anomalous scaling exponents q1 ; q4 ; gg10 ; gg11 , as well as scale invariance and statistical self-similarity arising from the competition between localization and delocalization gg12 ; q5 ; q6 . In paradigmatic settings such as the three-dimensional Anderson model, critical states occur at mobility edges separating extended and localized regimes NH3 ; Yao1 ; Yao2 .
Quasiperiodic systems cai ; Slager ; SSHQ ; gg3 ; Ayan1 ; Ayan2 provide an alternative platform for realizing critical states, even in the absence of true randomnes. Recent progress in critical states has uncovered a variety of nontrivial localization phenomena, including anomalous mobility edges Liu-Anomalous , renormalization-group descriptions RG1 ; RG2 , recursion-defined real spectra Liu-Real1 ; Liu-Real2 , and coupled-chain-induced criticality twochain . In previous theoretical descriptions Jiang1 ; Jiang2 ; gg5 ; gg6 , the asymptotic localization properties of single-particle eigenstates were typically characterized by a single real-space Lyapunov exponent $\gamma$. The Lyapunov exponent, as a dynamical indicator, signals the localization-delocalization transition. For exponentially localized states, the wave-function amplitudes decay as $|\psi_{n}|\propto e^{-\gamma|n-n_{0}|}$, where $n_{0}$ denotes the localization center and $\gamma>0$ quantifies the inverse localization length gg9 ; Duncan . A finite Lyapunov exponent $\gamma>0$ signals exponential localization, whereas $\gamma=0$ indicates delocalization. It is important to emphasize, however, that a vanishing Lyapunov exponent $\gamma=0$ in real space does not uniquely characterize extended states: both truly extended states and critical (multifractal) states fall into the non-exponentially localized class and exhibit delocalization. Consequently, a single-space description based solely on the real-space Lyapunov exponent suffers from intrinsic limitations when applied to critical states Liu-Anomalous .
Figure 1: (Color online) Schematic illustration of typical eigenstates in position and momentum space. Panels (a), (c), and (e) show extended, critical, and localized states in position space, respectively, while panels (b), (d), and (f) display the corresponding momentum-space distributions. From the dual-space perspective, critical states are uniquely distinguished by the absence of exponential localization in both spaces. This dual absence establishes the Lyapunov exponent as a duality invariant of multifractal critical states.
A useful conceptual perspective is provided by the Fourier uncertainty principle MQM , which dictates that localization in one representation necessarily implies delocalization in its Fourier-dual counterpart. Consequently, localized and extended states exhibit strongly asymmetric localization properties in position (real) and momentum spaces. This naturally raises the question of whether critical states, owing to their self-similar nature, may display a more symmetric behavior under position-momentum transformation.
Motivated by this consideration, Liu and Xia proposed a dual-space criterion for the exact identification of critical states Liu-Critical , overcoming the intrinsic limitations of single-representation descriptions. Specifically, they proposed that critical states lack exponential localization in both position and momentum space, exhibiting vanishing Lyapunov exponents and multifractal, nonergodic delocalization. By contrast, extended and localized states necessarily exhibit an inherent asymmetry between the two conjugate spaces: delocalization in one representation is accompanied by localization in the dual one, as illustrated in Fig. 1(a), (b), (e), and (f). Critical states, however, are uniquely characterized by invariant Lyapunov exponents, indicating delocalization in both position and momentum space, as shown in Fig. 1(c) and (d). This invariance of critical states under the Fourier transform, or dual-space transformation, is termed dual-space invariance.
Dual-space invariance of critical states was first introduced by Liu and Xia in an arXiv preprint in February 2023 Liu-Critical . Its applicability has since been supported in a variety of systems, including flat-band lattices quasi1 and Fibonacci topological insulators quasi2 . In March 2025, Ref. quasi3 claimed to characterize critical states through generalized incommensurate zeros in real and dual spaces, within the dual-space conceptual framework first proposed by Liu and Xia Liu-Critical . These developments underscore the growing interest in dual-space approaches to criticality and further highlight the foundational role of the dual-space invariance framework introduced by Liu et al., which predates and provides the conceptual basis for these subsequent studies.
The notion that critical states are delocalized in both position and momentum space provides an intuitive but oversimplified physical picture. This physical intuition is placed on a rigorous footing by the Liu–Xia criterion, which admits a precise mathematical formulation via the transfer-matrix method (see Appendix A). It identifies critical states through the simultaneous vanishing of the Lyapunov exponents in the two dual spaces,
$\gamma(E,V)=\gamma_{m}(E,V)=0,$ (1)
where $\gamma$ and $\gamma_{m}$ correspond to position and momentum space, respectively, $E$ is the eigenenergy, and $V$ the modulation strength. This implies that the Lyapunov exponent serves as a duality invariant of critical states, and its advantages become evident in comparison with finite-size multifractal analysis.
According to multifractal theory You ; quasi4 , in the thermodynamic limit, the minimal scaling index $\Gamma_{\min}$ approaches $1$ for extended states and 0 for localized states, while critical states satisfy $0<\Gamma_{\min}<1$. However, finite-size multifractal analysis suffers from intrinsic ambiguity, as critical states with small scaling exponents are difficult to distinguish from localized states. The Liu–Xia criterion overcomes this fundamental limitation by providing an exact and representation-independent identification of critical states. This is essential because the multifractal nature of critical states cannot be uniquely characterized by a single scaling exponent, as $\Gamma_{\min}$ spans a continuous interval $(0,1)$ rather than converging to a definite value. Finally, a unified framework for Anderson localization is summarized in Table 1.
Table 1: Classification of typical eigenstates in disordered quantum systems based on the Liu–Xia dual-space criterion Liu-Critical .
States Position-Momentum Delocalization Multifractal $\Gamma_{\min}$
Extended $\gamma=0,~\gamma_{m}>0$ Position Absent 1
Localized $\gamma>0,~\gamma_{m}=0$ Momentum Absent 0
Critical $\gamma=\gamma_{m}=0$ Both Present $(0,1)$
A natural question then arises: beyond Lyapunov exponents, does the dual-space characterization of critical states extend to other observables? Here, we show that experimentally accessible quantities, particularly the inverse participation ratio, exhibit analogous scaling in position and momentum space at criticality. This correspondence arises at the level of scaling and statistical properties rather than exact invariance. Our results therefore complement the Liu–Xia criterion and provide a practical route to identifying multifractal critical states in both theoretical and experimental settings.
II Phenomenological Analysis
Compared with the Lyapunov exponent, the inverse participation ratio (IPR) provides complementary information on wave-function localization Longhi1 ; Xia ; Jitomirskaya ; Avila ; IPR ; IPR1 ; IPR2 ; IPR3 . It is defined for normalized wave functions as
$\mathrm{IPR}=\sum_{n=1}^{L}|\psi_{n}|^{4},$ (2)
and quantifies the spatial confinement of the probability density. Extended states exhibit $\mathrm{IPR}\sim L^{-1}$ in the thermodynamic limit, whereas exponentially localized states are characterized by a finite IPR. Related quantities, such as the participation entropy, have also been widely used to probe localization and multifractality in theoretical and experimental studies Fan1 ; Fan2 ; Optical1 ; Optical2 ; gg0 ; gg1 ; gg2 .
For exponentially localized states, the finite IPR is consistent with a nonzero Lyapunov exponent. In contrast, critical states exhibit power-law decay and strong multifractal fluctuations, which preclude a description in terms of a single localization length. As a result, no exact analytical relation between the IPR and the Lyapunov exponent generally exists for critical states. Nevertheless, from a statistical perspective, the IPR is expected to display behavior analogous to that of the Lyapunov exponent.
Motivated by the Liu–Xia criterion Liu-Critical based on the Lyapunov exponent, we now investigate how the localization properties encoded in the IPR transform between position and momentum spaces. Although the IPR is inherently basis dependent and therefore not invariant under Fourier transformation (see the Appendix B), the dual-space invariance underlying the Liu–Xia criterion naturally suggests that critical states should exhibit the same scaling tendency in the two representations. This expectation can be formulated in an operational form as
$\mathrm{IPR}\sim\mathrm{IPR}_{m},$ (3)
where $\mathrm{IPR}_{m}$ denotes the inverse participation ratio evaluated in momentum space, and the symbol $\sim$” signifies an asymptotic correspondence at the level of scaling behavior, rather than a strict equality. Equation (3) should therefore be interpreted as a phenomenological signature of criticality, rather than an exact invariant under Fourier transformation.
By contrast, extended and exponentially localized states typically exhibit a pronounced asymmetry between position and momentum spaces, reflected in the markedly different scaling behavior of $\mathrm{IPR}$ and $\mathrm{IPR}_{m}$. In comparison, critical states remain protected by dual-space symmetry and exhibit comparable localization characteristics in both representations, consistent with their multifractal and self-similar nature. On this basis, Table 2 summarizes a phenomenological classification of eigenstates in disordered quantum systems.
Table 2: Phenomenological classification of eigenstates via the IPRs for finite-size systems.
States Position-Momentum Delocalization Multifractal $\Gamma_{\min}$
Extended $\rm IPR\sim 0,\rm IPR_{m}>0$ Position Absent 1
Localized $\rm IPR>0,\rm IPR_{m}\sim 0$ Momentum Absent 0
Critical $\rm IPR\sim\rm IPR_{m}$ Both Present $(0,1)$
This demonstrates that the Lyapunov exponent itself is an exact duality invariant for multifractal critical states, whereas the IPR, owing to its intrinsic basis dependence, is not strictly duality-invariant. It should be noted that the dual-space invariance discussed here refers to the invariance of localization characteristics at the level of scaling behavior and statistical properties, rather than an exact equality of critical wave functions. In systems with self-dual symmetry, the critical states at the self-dual point remain invariant under the position-momentum transformation. In contrast, without self-dual symmetry, a critical state in position space is mapped onto a distinct critical state in momentum space, while preserving the same scaling and statistical characteristics. This demonstrates that the dual-space invariance implied by the Liu–Xia criterion Liu-Critical is a genuinely nontrivial result. Rather than arising from self-duality that relates identical critical states, it reveals universal features shared by distinct critical states in position and momentum spaces, establishing a fundamentally new understanding of criticality in Anderson localization.
III Numerical verification
To verify previous predictions, we perform numerical simulations to investigate critical states in two representative quasiperiodic models. These systems are chosen because their Hamiltonians can be represented exactly in both position and momentum spaces, making them ideal for analysis.
III.1 Aubry-André-Harper model
The Aubry-André-Harper (AAH) model AAH1 ; AAH2 represents a fundamental paradigm for studying localization phenomena in quasiperiodic systems. In position space, its Hamiltonian takes the form of a discrete Schrödinger equation:
$\psi_{n+1}+\psi_{n-1}+V\cos(2\pi\alpha n+\theta)\psi_{n}=E\psi_{n}.$ (4)
where $\psi_{n}$ is the wave function amplitude of the particle at site $n$, $V_{n}=V\cos(2\pi\alpha n+\theta)$ is the quasiperiodic potential, and $E$ is the eigenvalue. $\alpha$ is an irrational number, usually taken to be the golden ratio $\alpha=\frac{{\sqrt{5}-1}}{2}$. $\theta$ is a phase factor, typically taken to be zero.
Utilizing the Fourier transform $\psi_{n}=\sum_{k}e^{-i2\pi\alpha nk}\phi_{k}$, the Hamiltonian of the AAH model in momentum space can be readily expressed,
$\frac{V}{2}(\phi_{k+1}+\phi_{k-1})+2\cos(2\pi\alpha k+\vartheta)\phi_{k}=E\phi_{k}.$ (5)
The AAH model exhibits a quantum phase transition driven by the quasiperiodic potential strength $V$. For weak quasiperiodic potential ($V<2$), all eigenstates are extended Bloch-like states, resulting in metallic transport. In contrast, for $V>2$, the system undergoes Anderson localization, where all eigenstates become exponentially localized, suppressing particle diffusion. At the critical point $V=2$, the Hamiltonians in real space [Eq. (4)] and momentum space [Eq. (5)] take identical forms. Consequently, the system reaches the self-dual point and exhibits multifractal eigenstates.
The procedure for identifying multifractal critical states according to the Liu–Xia criterion is outlined as follows. From Avila’s global theory Avila1 ; Avila2 , the Lyapunov exponents in position and momentum spaces are given by $\gamma=\ln(V/2)$ and $\gamma_{m}=\ln(2/V)$, respectively. Imposing the Liu–Xia criterion, $\gamma=\gamma_{m}$, rigorously and uniquely determines the critical point at $V=2$, thereby establishing the multifractal nature of the eigenstates. This result is in exact agreement with that obtained from the self-duality relation, further confirming the validity and consistency of the criterion.
III.2 Quasiperiodic-Nonlinear-Eigenproblem
The AAH model is well-known for its critical states at the self-dual point. So, does there exist a model without self-dual symmetry that still hosts critical states over a broad parameter range? Liu and Xia introduced such a model Liu-Critical , which is non-Hermitian. The construction of their model was inspired by the generalized eigenvalue problem. Mathematically, an eigenvalue problem takes the form:
$\hat{H}\left|\psi\right\rangle=E\hat{B}\left|\psi\right\rangle,$ (6)
where $\hat{H}$ denotes the Hamiltonian operator and $E$ represents the corresponding eigenvalue. The mathematical structure of Eq. (6) exhibits a fundamental dichotomy based on the form of $\hat{B}$. When $\hat{B}=I$, the equation reduces to a conventional linear eigenvalue problem. However, when $\hat{B}$ departs from the identity matrix, the problem transforms into either a nonlinear or generalized eigenvalue problem, introducing rich mathematical complexity absent in standard quantum systems.
Thus, the Quasiperiodic-Nonlinear-Eigenproblem (QNE) model Liu-Critical in momentum space can be formulated as
$\{2\cos[2\pi\alpha(k+1)]+V\}{\psi_{k+1}}+$ (7)
$\{2\cos[2\pi\alpha(k-1)]-V\}\psi_{k-1}=E(\psi_{k+1}+\psi_{k-1}).$
Non-Hermiticity in the QNE model originates from nonlinear eigenvalue terms, fundamentally distinct from the three established paradigms of asymmetric hopping, gain/loss, and complex potentials NH1 ; NH4 . This unconventional mechanism provides a new route to non-Hermitian spectra. As a direct consequence, the QNE model exhibits an expanded critical regime. Compared with the AAH model, where critical states exist only at the self-dual point $V=2$, the QNE model supports critical states over a broad parameter range $0<V\leq 2$.
After Fourier transformation, the real-space QNE Hamiltonian reduces to a linear eigenvalue problem that can be exactly diagonalized, dramatically reducing the computational cost (see Appendix C for details),
$\phi_{n+1}+\phi_{n-1}+iV\tan(2\pi\alpha n)\phi_{n}=E\phi_{n}.$ (8)
where $i$ represents the imaginary unit, and disregarding the phase factor.
Figure 2: (Color online) Panels (a1) and (a2) show the wave function of the AAH model ($V=2$, $E=0.0009$) in real and momentum space, respectively, while (b1) and (b2) show the corresponding results for the QNE model ($V=1$, $E=0.4$). The system size is $L=987$. Due to self-duality, critical states of the AAH model are invariant between real and momentum space. In contrast, the QNE model lacks self-duality: a critical state transforms into a distinct critical state under Fourier transformation, while their Lyapunov exponents remain identical.
Clearly, unlike the AAH model, the QNE model lacks self-duality symmetry, and its critical states generally exhibit different structures in real and momentum spaces. Consequently, the self-duality relation completely fails to identify the critical states. In sharp contrast, the Liu–Xia criterion remains fully operative. Specifically, the Lyapunov exponents in real and momentum spaces can be rigorously obtained from Eq. (8) and Eq. (7), respectively, and the critical states are uniquely determined by enforcing their equality, as demonstrated in Ref. Liu-Critical . Despite the absence of any self-duality symmetry, the critical states are still dictated by the invariance of the Lyapunov exponents in dual spaces, thereby revealing the fundamental origin, conceptual advance, and universal validity of the Liu–Xia criterion.
III.3 Verification results
Figure 3: (a) AAH model: MIPR and MIPRm exhibit distinct scaling away from the critical point $V=2$, but converge to comparable values at criticality. Results are averaged over all eigenstates for L=500, 1000, 2000L=500,\;1000,\;2000. (b) QNE model: in the regime $0<V<2$, both MIPR and MIPRm display strong fluctuations yet remain comparable, supporting criticality over a broadened parameter range. Results are obtained from six eigenstates near $E=0.00001$ for L=500, 1000, 2000L=500,\;1000,\;2000.
We perform direct numerical diagonalization for systems with $L=987$ to obtain the eigenvalues and eigenstates of both the AAH and QNE models. The results are shown in Fig. 2. As seen in panels (a1), (a2) and (b1), (b2), the critical states exhibit simultaneous delocalization, multifractality, and self-similarity in both position space ($|\psi|$) and momentum space ($|\phi|$). These results provide direct numerical evidence that critical wave functions obey dual-space invariance under Fourier transformation.
Figure 3 compares the mean inverse participation ratios in position space (MIPR) and momentum space (MIPRm) for the two quasiperiodic models. The MIPR is computed by averaging the IPR over eigenstates within a narrow energy window to suppress sample fluctuations. For the AAH model, the number of sampled states scales with system size, whereas for the QNE model, the averaging is limited to a few eigenstates due to computational cost.
As shown in Fig. 3(a), the AAH model exhibits strong asymmetry between position and momentum space away from criticality. For $V<2$, the real-space MIPR decreases with system size, consistent with extended states, while the momentum-space MIPRm remains finite, indicating localization in momentum space. The opposite behavior occurs for $V>2$. At the self-dual point $V=2$, the two quantities become comparable and follow similar scaling, reflecting dual-space invariance at criticality.
The behavior of the QNE model, shown in Fig. 3(b), is qualitatively different. For $V>2$, the eigenstates are exponentially localized in real space, leading to a weak size dependence of the MIPR, while the momentum-space MIPRm exhibits pronounced system-size scaling. In contrast, for $0<V<2$, both MIPR and MIPRm show strong fluctuations and nontrivial size dependence. Rather than approaching plateaus characteristic of extended or localized phases, the two quantities exhibit comparable scaling, consistent with a critical regime.
We emphasize that, due to finite-size effects and the basis dependence of the IPR, exact equality between MIPR and MIPRm is neither expected nor required. Instead, the results show that critical states are characterized by comparable scaling in position and momentum space. These findings provide direct numerical support for the dual-space correspondence.
IV Experimental feasibility
The dual-space correspondence identified in this work is well suited for experimental verification in ultracold atomic systems, where both position- and momentum-space observables are directly accessible. One-dimensional quasiperiodic lattices realizing AAH-type Hamiltonians can be implemented using optical lattices with superimposed incommensurate potentials.
In such setups, the real-space density distribution $|\psi_{n}|^{2}$ can be measured using in situ fluorescence or absorption imaging gg9 , allowing direct evaluation of the real-space IPR. The momentum-space distribution $|\phi_{k}|^{2}$ can be obtained via standard time-of-flight (TOF) measurements Biddle . After switching off the trapping and lattice potentials, the atomic cloud undergoes ballistic expansion, and in the long-TOF limit the measured density profile reflects the initial momentum distribution. By discretizing the TOF images according to the imaging resolution, the momentum-space inverse participation ratio $\mathrm{IPR}_{m}$ can be extracted in a directly analogous manner.
Within this framework, extended and localized states exhibit pronounced asymmetry between position and momentum space, whereas critical states show comparable localization properties in both representations. Although finite system sizes and experimental resolution prevent exact quantitative agreement, the scaling-level correspondence predicted here remains observable under realistic conditions. These results demonstrate that current cold-atom experiments provide a feasible and practical route for identifying multifractal critical states through dual-space diagnostics.
V Summary and outlook
In summary, we have investigated critical states in one-dimensional quasiperiodic systems from a dual-space perspective. Using analytical arguments and numerical simulations of representative models, including the AAH model and a quasiperiodic nonlinear eigenproblem, we have shown that critical states are uniquely characterized by comparable scaling in dual spaces. This property sharply distinguishes them from extended and localized states, which exhibit a pronounced asymmetry between position and momentum spaces.
We have further demonstrated that the inverse participation ratio provides a simple and practical diagnostic of this behavior: critical states exhibit comparable scaling of the inverse participation ratio in position and momentum space even in finite systems. This result establishes dual-space correspondence as an operational criterion for identifying multifractal criticality beyond conventional approaches based solely on a single space.
An important open direction is to examine interacting many-body systems, where quasiperiodicity and interactions can produce many-body localized and critical phases. It will be particularly interesting to determine whether a similar scaling correspondence emerges between real-space and Fock-space or momentum-space structures of many-body eigenstates, and whether the Liu–Xia criterion can be extended to provide a precise characterization of many-body criticality.
Acknowledgements.
This work was supported by the Natural Science Foundation of Nanjing University of Posts and Telecommunications (Grants No. NY223109).
Appendix A Computation of the Lyapunov exponents $\gamma$ and $\gamma_{m}$ via transfer-matrix methods
1. Real-space Lyapunov exponent $\gamma$
For a given energy $E$, the real-space Lyapunov exponent $\gamma(E)$ characterizes the asymptotic exponential growth of the wave-function amplitude along the lattice. It is computed using the standard transfer-matrix method. Starting from an initial vector $\mathbf{v}_{0}$ (e.g., $(1,0)^{\mathrm{T}}$), one iterates the site-dependent transfer matrix $\mathbf{T}_{n}(E)$,
$\mathbf{v}_{n+1}=\mathbf{T}_{n}(E)\mathbf{v}_{n}.$ (9)
After $N$ steps, the total transfer matrix is $\boldsymbol{\Pi}_{N}(E)=\mathbf{T}_{N-1}\cdots\mathbf{T}_{0}$, and the Lyapunov exponent is defined as
$\gamma(E)=\lim_{N\to\infty}\frac{1}{N}\ln\left\|\boldsymbol{\Pi}_{N}(E)\right\|.$ (10)
In numerical implementation, periodic renormalization is applied to prevent overflow, and $\gamma(E)$ is obtained from the averaged logarithmic growth rate for large $N$ (typically $N\sim 10^{5}$).
2. Momentum-space Lyapunov exponent $\gamma_{m}$
The momentum-space Lyapunov exponent $\gamma_{m}(E)$ is computed using an analogous transfer-matrix formulation in momentum space. After Fourier transforming the eigenvalue equation, the momentum-space amplitudes $\tilde{\psi}_{k}$ satisfy a linear recurrence relation. For quasiperiodic potentials with a single incommensurate frequency, this relation forms an effective one-dimensional chain,
$a_{k}\tilde{\psi}_{k+Q}+b_{k}\tilde{\psi}_{k}+c_{k}\tilde{\psi}_{k-Q}=E\tilde{\psi}_{k},$ (11)
which can be rewritten as
$\begin{pmatrix}\tilde{\psi}_{k+Q}\\ \tilde{\psi}_{k}\end{pmatrix}=\mathbf{M}_{k}(E)\begin{pmatrix}\tilde{\psi}_{k}\\ \tilde{\psi}_{k-Q}\end{pmatrix}$. (12)
Iterating the corresponding transfer matrices yields
$\gamma_{m}(E)=\lim_{K\to\infty}\frac{1}{K}\ln\left\|\boldsymbol{\Pi}^{(m)}_{K}(E)\right\|,$ (13)
where $\boldsymbol{\Pi}^{(m)}_{K}(E)$ denotes the product of momentum-space transfer matrices.
In both cases, the Lyapunov exponent quantifies the exponential growth rate of the transfer-matrix norm, with the only difference being whether the recurrence is defined in position or momentum space.
Appendix B Contrast Between Basis-Dependent IPR and Basis-Invariant Lyapunov Exponents
It is instructive to contrast the distinct roles of the Lyapunov exponent and the inverse participation ratio (IPR) in characterizing localization. The Lyapunov exponent is defined by the asymptotic exponential growth or decay of solutions to a linear recurrence relation and thus represents an intrinsic dynamical quantity associated with the transfer-matrix structure. Consequently, it is independent of the choice of basis and depends only on the underlying recursive dynamics in a given representation (real or momentum space).
In contrast, the IPR depends explicitly on the expansion coefficients of the wave function in a chosen basis and quantifies the degree of localization within that representation. Because higher-order moments of the wave-function amplitudes are not invariant under unitary transformations, the IPR is inherently basis dependent and generally differs between position and momentum space.
This distinction is central to the present work. While the simultaneous vanishing of Lyapunov exponents in dual spaces provides a basis-invariant criterion for criticality, the emergence of comparable scaling of the IPR in dual spaces constitutes a nontrivial physical signature. It indicates that, despite its basis dependence, the IPR reflects the reduced asymmetry between position and momentum space at criticality.
Appendix C Fourier transformation for the QNE model
We show that the real-space equation
$\phi_{n+1}+\phi_{n-1}+iV\tan(2\pi\alpha n)\phi_{n}=E\phi_{n}$ (14)
transforms into the momentum-space equation
$\bigl[2\cos\bigl(2\pi\alpha(k+1)\bigr)+V\bigr]\psi_{k+1}$ (15)
$\quad+\bigl[2\cos\bigl(2\pi\alpha(k-1)\bigr)-V\bigr]\psi_{k-1}=E(\psi_{k+1}+\psi_{k-1}),$
under a discrete Fourier transform.
1. Fourier transform
Multiplying Eq. (14) by $\cos(2\pi\alpha n)$ gives
$\cos(2\pi\alpha n)\phi_{n+1}+\cos(2\pi\alpha n)\phi_{n-1}$ (16)
$\quad+iV\sin(2\pi\alpha n)\phi_{n}=E\cos(2\pi\alpha n)\phi_{n}.$
We define the discrete Fourier transform
$\psi_{k}=\sum_{n}e^{-i2\pi\alpha kn}\phi_{n},\qquad\phi_{n}=\sum_{k}e^{i2\pi\alpha kn}\psi_{k}.$ (17)
2. Momentum-space equation
Using standard trigonometric identities, the Fourier transform yields
$\sum_{n}e^{-i2\pi\alpha kn}\cos(2\pi\alpha n)\phi_{n\pm 1}$ $=\cos\bigl[2\pi\alpha(k\pm 1)\bigr]\psi_{k\pm 1},$ (18)
$\sum_{n}e^{-i2\pi\alpha kn}\sin(2\pi\alpha n)\phi_{n}$ $=\frac{1}{2i}\bigl(\psi_{k-1}-\psi_{k+1}\bigr),$ (19)
$\sum_{n}e^{-i2\pi\alpha kn}\cos(2\pi\alpha n)\phi_{n}$ $=\frac{1}{2}\bigl(\psi_{k-1}+\psi_{k+1}\bigr).$ (20)
Substituting into Eq. (16) gives
$\cos\bigl[2\pi\alpha(k+1)\bigr]\psi_{k+1}+\cos\bigl[2\pi\alpha(k-1)\bigr]\psi_{k-1}$ (21)
$\quad+\frac{V}{2}(\psi_{k-1}-\psi_{k+1})=\frac{E}{2}(\psi_{k-1}+\psi_{k+1}),$
which reduces to Eq. (15).